OpenStudy (anonymous):

Verify that cot x sec4x = cot x + 2 tan x + tan3x

OpenStudy (anonymous):

so i know cotx is cos/sin

OpenStudy (anonymous):

idk if we can change the tan though to sin/cos?

OpenStudy (anonymous):

this one is hard but try this

OpenStudy (anonymous):

first divide everything by cot x

OpenStudy (anonymous):

\[\sec^{4} x = 1 + 2\frac{ \tan x }{ \cot x } + \frac{ \tan^{3} x }{ \cot x }\]

OpenStudy (anonymous):

and we know that \[\frac{ \tan x }{ \cot x } = \frac{ \frac{ \sin x }{ \cos x } }{ \frac{ \cos x }{ \sin x } } = \frac{\sin^{2} x }{ \cos^{2} x } = \tan^{2} x\]

OpenStudy (anonymous):

would it apply to the second one but it would be (sin^3x/cos^3x)/cosx/sinx ?

OpenStudy (anonymous):

that would simplify it to: \[\sec^{4} x = 1 + 2 \tan^{2} x + \tan^{4} x\]

OpenStudy (anonymous):

yeah the second one is: \[\frac{ \tan^{3} x }{ \cot x } = \frac{ \tan x }{ \cot x } * \tan^{2} x = \tan^{2} x * \tan^{2} x = \tan^{4} x\]

OpenStudy (anonymous):

where did you get the second tan^2x

OpenStudy (anonymous):

I divided tan^3 x into tan^x * tan^2 x

OpenStudy (anonymous):